# Wukong’s 72 Transformations: High-fidelity Textured 3D Morphing via Flow Models

### v3

@@ Abstract
We present Wukong, a novel training-free framework for high-fidelity textured 3D morphing that takes a pair of source and target prompts (image or text) as input. Unlike conventional methods—which rely on manual correspondence matching and deformation trajectory estimation (limiting generalization and requiring costly preprocessing)—Wukong leverages the generative prior of flow-based transformers to produce high-fidelity 3D transitions with rich texture details. To ensure smooth shape transitions, we exploit the inherent continuity of flow-based generative processes and formulate morphing as an optimal transport barycenter problem. We further introduce a sequential initialization strategy to prevent abrupt geometric distortions and preserve identity coherence. For faithful texture preservation, we propose a similarity-guided semantic consistency mechanism that selectively retains high-frequency details and enables precise control over blending dynamics. This avoids common artifacts like oversmoothing while maintaining semantic fidelity. Extensive quantitative and qualitative evaluations demonstrate that Wukong significantly outperforms state-of-the-art methods, achieving superior results across diverse geometry and texture variations. Project page: https://visual-ai.github.io/wukong
@@ 1 Introduction
Recently, 2D image morphing has achieved remarkable success (qiyuan2024aid; zhang2023diffmorpher), driven by advances in 2D diffusion models (e.g., Stable Diffusion (rombach2022high)). Inspired by this progress, we aim to extend these successes to 3D – morphing not only shapes but also textures – without relying on large-scale paired data. In other words, we aim to develop a training-free framework for textured 3D morphing by leveraging the strong priors of generative models. However, building such a framework poses significant challenges due to the absence of large-scale 3D data with continues shape and texture transitions.
In this paper, we propose Wukong, a novel framework for high-fidelity textured 3D morphing that takes image or text pairs as input to delineate source and target objects. Built upon a pre-trained flow-based transformer (xiang2024structured) for 3D generation, Wukong bridges the gap between 3D shapes and image/text conditions. Leveraging the deterministic property of flow models (where intermediate states are not stochastic), we derive morphing trajectories by solving a free-support Wasserstein barycenter problem. Additionally, we introduce a sequential initialization strategy to enhance the smoothness of the transitions along the barycentric trajectory. This design also allows us to handle the texture morphing in 3D in a similar fashion. However, this is not sufficient to produce faithful texture. As naïve interpolation leads to undesirable blending artifacts (e.g., loss of facial details and color patterns). To address this, we propose a similarity-guided consistency mechanism that selectively preserves high-frequency texture details while providing finer control over transition dynamics.
The main contributions of this work are as follows: (i) We propose Wukong, a novel framework for high-fidelity textured 3D morphing that takes a pair of image or text prompts as input. By leveraging a flow-based generative model as a prior, Wukong enables smooth and controllable shape and texture interpolation. (ii) We introduce a method to derive faithful intermediate morphing states by solving an optimal transport barycenter problem, further augmented by a sequential initialization strategy to facilitate smooth transitions. (iii) To address texture degradation in morphing, we propose a similarity-guided semantic consistency mechanism that selectively preserves high-frequency texture details, enabling finer control over texture transitions. (iv) Through extensive experiments on diverse 3D morphing scenarios—spanning different object categories with varying geometry and significant texture changes—we demonstrate that Wukong outperforms existing methods, establishing a new state-of-the-art in textured 3D morphing.
@@ 3.1 Overview and formulation
Figure 2: Our Wukong framework. Given a source and a target (image or text), we extract features using pretrained encoders and treat the condition tokens as empirical distributions. We compute their Wasserstein barycenter with weights $(1-\beta)$ and $\beta$ to obtain interpolated condition tokens $\mathcal{Z}$. These are fed into a shared geometry flow model $\Phi_{geo}$ and texture flow model $\Phi_{tex}$ to generate 3D outputs at different $\alpha$ values, producing textured 3D morphs. The top-right shows our texture-controlled morphing branch, and the bottom-right illustrates the recursive initialization strategy.
Given two input prompts ($P_{\text{source}}$, $P_{\text{target}}$) in image or text form, our method learns an interpolation function $\mathbf{I}$ to generates a coherent sequence of 3D textured meshes $\mathcal{G}=\{\mathbf{G}_{\alpha}\}_{\alpha=0}^{J+1}$, which forms a smooth morphing trajectory between the two concepts. Here, $\alpha\in\{0,\ldots,J+1\}$ denotes the discrete interpolation step index. And $\mathbf{G}_{0}$ and $\mathbf{G}_{J+1}$ denote the 3D generation of the input prompts $P_{\text{source}}$ and $P_{\text{target}}$, respectively. Formally, the problem can be formulated as:
@@ 3.2 Flow-based 3D generator
(1) Geometric structure generation that generates sparse structure $\boldsymbol{p}=\{p_{i}\}_{i=1}^{L}$ from the condition $\mathcal{C}$, which is either text-embedded CLIP features (radford2021learning) or image-encoded DINOv2 features (oquab2023dinov2):
@@ 3.4 Controllable texture generation via selective interpolation
Although our shape-level interpolation generates semantically coherent 3D transitions, it tends to produce undesirable texture averaging, particularly for fine details like facial features, color patterns, and garment elements, due to its uniform application across all token regions. To overcome this limitation, we introduce a selective interpolation strategy that: (1) preserves high-frequency texture details through similarity-guided blending, while (2) maintaining smooth semantic transitions at the structural level. This approach enables controlled transfer of local texture attributes from one condition while harmoniously integrating global characteristics from the other.
Building upon the geometric interpolation introduced in Sec. 3.3, we first compute the free-support barycenter $\mathcal{Z}_{\alpha}=\{\mathbf{z}_{k}\}_{k=1}^{n}$ from Eq. 9. While this provides a general interpolation baseline, it may dilute fine-grained details from the source. To mitigate this, we augment the interpolation with a semantic consistency evaluation between interpolated points $\mathbf{z}_{k}$ and the source tokens $\mathcal{X}$, $\mathcal{Y}$.
@@ 4 Experimental results
We adopt the Trellis framework as our 3D generative model. Both the structure flow and texture flow models are implemented using rectified flow with 25 sampling steps each. The classifier-free guidance (CFG) strength is set to 3. DINOv2 (oquab2023dinov2) and CLIP (radford2021learning) are used for image and text feature extraction. Both the structure and texture flow-based generator $\Phi_{\text{geo}}$, $\Phi_{\text{tex}}$ contain 21 cross-attention blocks, where interpolation is performed on every condition token before each cross-attention layer. The morphing coefficient $\alpha$ is uniformly sampled during morphing and we set $J=6$ for experiments presented in the paper. For shape interpolation, we compute the token-wise barycenter using the free-support Wasserstein barycenter implemented by ot.lp.free_support_barycenter (lindheim2023simple). We set the maximum number of optimization iterations to 100, and the convergence threshold (stop criterion) to $1\times 10^{-5}$. We conduct experiments on an NVIDIA A100. Generating a single morphed 3D output takes approximately 30 seconds.
@@ 4.1.1 Quantitative results
In Table LABEL:tab:quantitative, we compare our method against a range of baselines across multiple evaluation metrics. As can be observed, our method consistently outperforms existing approaches across all metrics, demonstrating superior quality and consistency in both shape and texture morphing. Specifically, we outperform state-of-the-art textured 3D morphing methods across the board, i.e., 3DRM (yang2025textured) and MorphFlow (tsai2022multiview), showing clear improvements in perceptual quality and semantic coherence. Furthermore, compared with image-based morphing methods (zhang2023diffmorpher; qiyuan2024aid; jin2024mv; luma2025dreammachine), our approach also exhibits much stronger 3D consistency and higher fidelity in both appearance and structure. See Appendix B for more quantitative evaluations.
Figure 3: 3D morphing with image prompts. The leftmost column shows source image prompts, while the rightmost column shows target prompts. Intermediate columns depict the morphing trajectory generated by our method.
Figure 4: 3D morphing with text prompts. The leftmost column shows the source text descriptions, and the rightmost column shows the target prompts. Intermediate results visualize the smooth morphing trajectory generated by our method.
@@ 4.1.2 Qualitative results
Figure 6: Shape and texture interpolation results. The figure shows shape interpolation arranged horizontally, with each row representing a morphing process at different interpolation thresholds, illustrating the smooth transition from the source image to the target image.
@@
In Fig. 7, we demonstrate the controllability of Wukong in textured 3D morphing through our proposed selective texture interpolation. Each row represents a distinct morphing process with different interpolation thresholds, with the leftmost column fixed as the source shape. The results along the vertical axis (columns) are independent of each other. The far-left column does not show texture changes, as it represents the same 3D source generated from the input source image. By adjusting the selective interpolation threshold $\tau$ (from top to bottom), we can control the relative influence of source and target tokens during the morphing process. For instance, we can preserve source-dominant features such as facial identity, clothing styles, and color patterns in the intermediate 3D shapes, while still achieving smooth and semantically coherent shape transitions.
@@ 4.2 Ablation study
We further conduct ablation studies for shape interpolation under the setting of text-conditioned 3D morphing, as in Fig. 8. The results reveal the following: (1) Linear interpolation (row 1) produces unnatural artifacts, such as T-pose human figures, indicating that direct token blending often strays into semantically invalid regions. (2) Our method without recursive initialization (row 2) exhibits similar issues, underscoring the importance of proper initialization for barycenter optimization. Linear initialization fails to respect the underlying data manifold, often leading to unstable or artifact-prone results, as detailed in Sec. 3.3. (3) Our full method (row 3), with recursive initialization, successfully avoids these problems, delivering smooth and structural plausible shape transitions throughout the morph.
@@
For quantatitive comparison, Table LABEL:tab:ablation provides further ablation results to validate the effectiveness of our method. The first row (“Linear”) uses direct linear interpolation between latent tokens without any optimization or structural guidance. This naive baseline leads to unnatural intermediate shapes and visible texture artifacts. The second row (“w/o r-i”) uses our barycenter-based optimization but removes the recursive initialization procedure. Although this improves upon linear interpolation by refining support points through optimization, the lack of a good initialization often causes the optimization to converge to suboptimal solutions, resulting in distortions and unstable transitions. The third row (“Ours*”) corresponds to our full method applied on the TripoSG (li2025triposg) backbone. This configuration includes our recursive initialization scheme, which provides a strong prior for the barycenter and guides the optimization to more stable and meaningful intermediate representations. Compared to prior rows, “Ours*” produces smoother geometry transitions and better texture blending, even in challenging cases such as cross-category morphing. Notably, both “Ours*” and our default backbone implementation (“Ours”) achieve the highest visual fidelity and semantic consistency across the morphing trajectory, demonstrating that our method generalizes well across different 3D generative models and significantly improves interpolation quality through careful design of both initialization and optimization procedures.
@@
In Fig. 7 (rows 4 &\& 5), we demonstrate the importance of our selective interpolation strategy. In row 4, directly increasing source token weights via naïve linear interpolation leads to texture collapse, with disorganized colors and unclear semantics. In contrast, row 5 shows that our selective interpolation yields smooth morphing results while preserving distinctive source features like Wukong’s appearance and attire. This demonstrates that selective control is crucial for high-quality and identity-preserving texture transitions.
We conduct quantitative ablation experiments for texture interpolation strategy as shown in Table LABEL:tab:ablation_texture_appendix. The first row “w/o interp” refers to directly copying the source texture tokens across all steps without any interpolation. The second row “linear” refers to applying standard linear interpolation between source and target texture tokens. The third row “ours” corresponds to using our proposed selective interpolation strategy based on similarity thresholding. Our method achieve the highest visual fidelity and semantic consistency throughout the morphing trajectory, demonstrating that our method generalizes effectively across various 3D generative models. Additionally, it significantly enhances interpolation quality by carefully designing both the texture interpolation procedures.
@@ Appendix A Model details
To enable high-fidelity textured 3D morphing, we build upon two pretrained flow-based transformer models introduced in Trellis (xiang2024structured): structure flow model and SLat flow model, originally designed for unconditional 3D generation. The structure flow model operates on structured latent representation and follows a transformer-based architecture with 24 modulated transformer blocks with cross attentions. Each block contains self-attention, cross-attention, and feed-forward components, modulated via AdaLN (guo2022adaln) conditioning from a learned timestep embedding. Root Mean Square Normalization (RMSNorm) (zhang2019root) is applied to both the query and key representations prior to their use in the attention mechanism. The SLat flow model incorporates a hierarchical design with sparse 3D convolutional blocks and positional embeddings to encode spatial context. The transformer core comprises 24 modulated sparse transformer blocks with cross attentions, analogous in structure to the geometry model but enhanced with sparse attention and feed-forward operations. Additionally, the model includes dedicated output convolutional blocks for upsampling and decoding, ensuring fine-grained preservation and modulation of high-frequency texture details. We use the free-support Wasserstein barycenter solver from the POT library (ot.lp.free_support_barycenter (lindheim2023simple)), which is based on linear programming (LP). The cost matrix is computed using the squared Euclidean distance in the CLIP (radford2021learning) (for text) and DINOv2 (oquab2023dinov2) (for image) embedding space.

### v4

@@ Abstract
We present Wukong, a novel training-free framework for high-fidelity textured 3D morphing that takes a pair of source and target prompts (image or text) as input. Unlike conventional methods—which rely on manual correspondence matching and deformation trajectory estimation (limiting generalization and requiring costly preprocessing)—Wukong leverages the generative prior of flow-based transformers to produce high-fidelity 3D transitions with rich texture details. To ensure smooth shape transitions, we exploit the inherent continuity of flow-based generative processes and formulate morphing as an optimal transport barycenter problem. We further introduce a sequential initialization strategy to prevent abrupt geometric distortions and preserve identity coherence. For faithful texture preservation, we propose a similarity-guided semantic consistency mechanism that selectively retains high-frequency details and enables precise control over blending dynamics. This empowers Wukong to support both global texture transitions and identity-preserving texture morphing, catering to diverse generation needs. Extensive quantitative and qualitative evaluations demonstrate that Wukong significantly outperforms state-of-the-art methods, achieving superior results across diverse geometry and texture variations. Project page: https://visual-ai.github.io/wukong
@@ 1 Introduction
Recently, 2D image morphing has achieved remarkable success (qiyuan2024aid; zhang2023diffmorpher), driven by advances in 2D diffusion models (e.g., Stable Diffusion (rombach2022high)). Inspired by this progress, we aim to extend these successes to 3D – morphing not only shapes but also textures – without relying on large-scale paired data. In other words, we aim to develop a training-free framework for textured 3D morphing by leveraging the strong priors of generative models. However, building such a framework poses significant challenges due to the absence of large-scale 3D data with continuous shape and texture transitions.
In this paper, we propose Wukong, a novel framework for high-fidelity textured 3D morphing that takes image or text pairs as input to delineate source and target objects. Built upon a pre-trained flow-based transformer (xiang2024structured) for 3D generation, Wukong bridges the gap between 3D shapes and image/text conditions. Leveraging the deterministic property of flow models (where intermediate states are not stochastic), we derive morphing trajectories by solving a free-support Wasserstein barycenter problem. Additionally, we introduce a sequential initialization strategy to enhance the smoothness of the transitions along the barycentric trajectory. This design also allows us to handle the texture morphing in 3D in a similar fashion. However, a naive interpolation treats all texture regions uniformly, which may not satisfy the need for preserving specific source attributes, such as fine details, distinct styles, or identity patterns. To address this, we propose a similarity-guided consistency mechanism for texture controlled morphing, which selectively preserves high-frequency texture details while providing finer control over transition dynamics.
The main contributions of this work are as follows: (i) We propose Wukong, a novel framework for high-fidelity textured 3D morphing that takes a pair of image or text prompts as input. By leveraging a flow-based generative model as a prior, Wukong enables smooth and controllable shape and texture interpolation. (ii) We introduce a method to derive faithful intermediate morphing states by solving an optimal transport barycenter problem, further augmented by a sequential initialization strategy to facilitate smooth transitions. (iii) We propose a similarity-guided semantic consistency mechanism to enable texture controlled morphing, allowing users to selectively preserve high-frequency texture details and exercise fine-grained control over appearance transitions. (iv) Through extensive experiments on diverse 3D morphing scenarios—spanning different object categories with varying geometry and significant texture changes—we demonstrate that Wukong outperforms existing methods, establishing a new state-of-the-art in textured 3D morphing.
@@ 3.1 Overview and formulation
Figure 2: Our Wukong framework. Given a source and a target (image or text), we extract features using pretrained encoders and treat the condition tokens as empirical distributions. We compute their Wasserstein barycenter with weights $(1-\beta)$ and $\beta$ to obtain interpolated condition tokens $\mathcal{Z}$. These tokens are fed into a shared geometry flow model $\Phi_{geo}$ and texture flow model $\Phi_{tex}$ to generate 3D outputs at different $\alpha$ values, producing textured 3D morphs. The top-right shows our texture-controlled morphing branch, and the bottom-right illustrates the recursive initialization strategy.
Given two input prompts ($P_{\text{source}}$, $P_{\text{target}}$) in image or text form, our method learns an interpolation function $\mathbf{I}$ to generate a coherent sequence of 3D textured meshes $\mathcal{G}=\{\mathbf{G}_{\alpha}\}_{\alpha=0}^{J+1}$, which forms a smooth morphing trajectory between the two concepts. Here, $\alpha\in\{0,\ldots,J+1\}$ denotes the discrete interpolation step index. Among them, $\mathbf{G}_{0}$ and $\mathbf{G}_{J+1}$ denote the 3D generation of the input prompts $P_{\text{source}}$ and $P_{\text{target}}$, respectively. Formally, the problem can be formulated as:
@@ 3.2 Flow-based 3D generator
(1) Geometric structure generation that generates sparse structure $\boldsymbol{p}=\{p_{i}\}_{i=1}^{L}$ from the condition $\mathcal{C}$, which are either text-embedded CLIP features (radford2021learning) or image-encoded DINOv2 features (oquab2023dinov2):
@@ 3.4 Controllable texture generation via selective interpolation
As illustrated in Fig. 2, our framework supports two strategies for texture evolution. The default textured 3D morphing utilizes the barycenter interpolated tokens $\mathcal{Z}_{\alpha}$ (derived in Sec. 3.3) as the condition for the texture flow model $\Phi_{tex}$. Although this produces a smooth transformation, it strictly couples texture evolution with the structural transition. However, creative applications often demand more flexibility—specifically, the ability to preserve the source’s visual identity (e.g., facial features or iconic patterns) even as the shape transforms. One may consider enforcing source tokens $\mathcal{X}$ across all token regions, but this is not feasible: as the geometry deforms, forcing static tokens onto the evolving structure creates inherent texture-geometry misalignment, resulting in visual artifacts.
To address this, we introduce texture controlled 3D morphing. This mode employs a selective interpolation strategy that functions as a tunable filter. By adjusting a similarity threshold, users can explicitly regulate the degree of source preservation, spanning from subtle detail retention to strong identity maintenance. This mechanism retains high-frequency details from $\mathcal{X}$ only where semantically aligned to prevent artifacts, while utilizing $\mathcal{Z}_{\alpha}$ elsewhere to ensure structural coherence.
Building upon the geometric interpolation introduced in Sec. 3.3, we first compute the free-support barycenter $\mathcal{Z}_{\alpha}=\{\mathbf{z}_{k}\}_{k=1}^{n}$ from Eq. 9. While this provides a general interpolation baseline, it may dilute fine-grained details from the source. To mitigate this, we augment the interpolation with a semantic consistency evaluation between barycenter points $\mathbf{z}_{k}$ and the source tokens $\mathcal{X}$, $\mathcal{Y}$.
@@ 4 Experimental results
We adopt the Trellis framework as our 3D generative model. Both the structure flow and texture flow models are implemented using rectified flow with 25 sampling steps each. The Classifier-Free Guidance (CFG) strength is set to 3. DINOv2 (oquab2023dinov2) and CLIP (radford2021learning) are used for image and text feature extraction. Both the structure and texture flow-based generator $\Phi_{\text{geo}}$, $\Phi_{\text{tex}}$ contain 21 cross-attention blocks, where interpolation is performed on every condition token before each cross-attention layer. The morphing coefficient $\alpha$ is uniformly sampled during morphing and we set $J=6$ for experiments presented in the paper. For shape interpolation, we compute the token-wise barycenter using the free-support Wasserstein barycenter implemented by ot.lp.free_support_barycenter (lindheim2023simple). We set the maximum number of optimization iterations to 100, and the convergence threshold (stop criterion) to $1\times 10^{-5}$. We conduct experiments on an NVIDIA A100 GPU. Generating a single morphed 3D output takes approximately 30 seconds.
@@ 4.1.1 Quantitative results
In Tab LABEL:tab:quantitative, we compare our method against a range of baselines across multiple evaluation metrics. As can be observed, our method consistently outperforms existing approaches across all metrics, demonstrating superior quality and consistency in both shape and texture morphing. Specifically, we outperform state-of-the-art textured 3D morphing methods across the board, i.e., 3DRM (yang2025textured) and MorphFlow (tsai2022multiview), showing clear improvements in perceptual quality and semantic coherence. Furthermore, compared with image-based morphing methods (zhang2023diffmorpher; qiyuan2024aid; jin2024mv; luma2025dreammachine), our approach also exhibits much stronger 3D consistency and higher fidelity in both appearance and structure. See Appendix B for more quantitative evaluations.
Figure 3: Textured 3D morphing with image prompts. The leftmost column shows source image prompts, while the rightmost column shows target prompts. Intermediate columns depict the morphing trajectory generated by our method.
Figure 4: Textured 3D morphing with text prompts. The leftmost column shows the source text descriptions, and the rightmost column shows the target prompts. Intermediate results visualize the smooth morphing trajectory generated by our method.
@@ 4.1.2 Qualitative results
Figure 6: Texture controlled morphing results. The figure shows shape interpolation arranged horizontally, with each row representing a morphing process at different interpolation thresholds, illustrating the smooth transition from the source image to the target image.
@@
In Fig. 7, we demonstrate the capabilities of our texture controlled morphing mode, by employing our proposed selective texture interpolation. Each row represents a distinct morphing process with different interpolation thresholds, with the leftmost column fixed as the source shape. The results along the vertical axis (columns) are independent of each other. The far-left column does not show texture changes, as it represents the same 3D source generated from the input source image. By adjusting the selective interpolation threshold $\tau$ (from top to bottom), we can control the relative influence of source and target tokens during the morphing process. For instance, we can preserve source-dominant features such as facial identity, clothing styles, and color patterns in the intermediate 3D shapes, while still achieving smooth and semantically coherent shape transitions.
@@ 4.2 Ablation study
We further conduct ablation studies for shape interpolation under the setting of text-conditioned 3D morphing, as in Fig. 8. The results reveal the following: (1) Linear interpolation (row 1) produces unnatural artifacts, such as T-pose human figures, indicating that direct token blending often strays into semantically invalid regions. (2) Our method without recursive initialization (row 2) exhibits similar issues, underscoring the importance of proper initialization for barycenter optimization. Linear initialization fails to respect the underlying data manifold, often leading to unstable or artifact-prone results, as detailed in Sec. 3.3. (3) Our full method (row 3), with recursive initialization, successfully avoids these problems, delivering smooth and structurally plausible shape transitions throughout the morph.
@@
For quantitative comparison, Tab LABEL:tab:ablation provides ablation results to validate our method. The “Linear” baseline, using direct interpolation without optimization, leads to unnatural shapes and visible artifacts. The “w/o r-i” row employs optimization but omits recursive initialization; while it improves upon the linear baseline, the lack of a robust prior often causes convergence to suboptimal solutions, resulting in distortions. In contrast, “Ours*” applies our full method to the TripoSG (li2025triposg) backbone. This configuration includes recursive initialization, which acts as a strong prior to guide optimization toward stable and meaningful representations. Consequently, “Ours*” yields smoother geometry and better texture blending than the baselines. Notably, both “Ours*” and our default “Ours” achieve the highest fidelity, demonstrating that our method generalizes well across models and improves quality through effective initialization and optimization.
@@
In Fig. 7 (rows 4 &\& 5), we evaluate the effectiveness of texture controlled morphing enabled by our selective interpolation strategy. In row 4, directly increasing source token weights via naïve linear interpolation leads to texture collapse, with disorganized colors and unclear semantics. In contrast, row 5 shows that our selective interpolation yields smooth morphing results while preserving distinctive source features like Wukong’s appearance and attire. This demonstrates that selective control is crucial for high-quality and identity-preserving texture transitions.
We conduct quantitative ablation experiments for texture interpolation strategy as shown in Tab LABEL:tab:ablation_texture_appendix. The first row “w/o interp” refers to directly copying the source texture tokens across all steps without any interpolation. The second row “linear” refers to applying standard linear interpolation between source and target texture tokens. The third row “ours” corresponds to using our proposed selective interpolation strategy based on similarity thresholding. Our method achieves the highest visual fidelity and semantic consistency throughout the morphing trajectory, demonstrating that our method generalizes effectively across various 3D generative models. Additionally, it significantly enhances interpolation quality by carefully designing both the texture interpolation procedures.
@@ Appendix A Model details
To enable high-fidelity textured 3D morphing, we build upon two pretrained flow-based transformer models introduced in Trellis (xiang2024structured): the structure flow model and SLat flow model, originally designed for unconditional 3D generation. The structure flow model operates on a structured latent representation and follows a transformer-based architecture with 24 modulated transformer blocks with cross attentions. Each block contains self-attention, cross-attention, and feed-forward components, modulated via AdaLN (guo2022adaln) conditioning from a learned timestep embedding. Root Mean Square Normalization (RMSNorm) (zhang2019root) is applied to both the query and key representations prior to their use in the attention mechanism. The SLat flow model incorporates a hierarchical design with sparse 3D convolutional blocks and positional embeddings to encode spatial context. The transformer core comprises 24 modulated sparse transformer blocks with cross attentions, analogous in structure to the geometry model but enhanced with sparse attention and feed-forward operations. Additionally, the model includes dedicated output convolutional blocks for upsampling and decoding, ensuring fine-grained preservation and modulation of high-frequency texture details. We use the free-support Wasserstein barycenter solver from the POT library (ot.lp.free_support_barycenter (lindheim2023simple)), which is based on linear programming (LP). The cost matrix is computed using the squared Euclidean distance in the CLIP (radford2021learning) (for text) and DINOv2 (oquab2023dinov2) (for image) embedding space.